OpenStudy (anonymous):

solve for:4cos(pi)x+5=7

OpenStudy (anonymous):

try this subtract 5 both sides 4cos(pi)x=2 cos(pi)x=(1/2) cos inverse both sides (pi)x=60 degree 60 degree also called (pi/3) so final answer is x=1/3 may be i am wrong

OpenStudy (anonymous):

u r correct

OpenStudy (anonymous):

Thanks jimmy

OpenStudy (anonymous):

yw